OpenStudy (anonymous):

Quadratic equations????? I tried to solve it but I came out with something completely different. Solve 6x2 = 12x − 24. the quantity of negative 2 plus or minus 3i square root 2 all over 2 1 plus or minus i square root 3 the quantity of 2 plus or minus 3i square root 2 all over 2 1 plus or minus 2i square root 10

OpenStudy (anonymous):

All terms have a common factor of 6, so divide all terms by 6.

Vocaloid (vocaloid):

start by rewriting it in the form ax^2 + bx + c = 0 6x^2 - 12x + 24 = 0 then apply the quadratic formula

Vocaloid (vocaloid):

(you can factor out 6)

OpenStudy (anonymous):

Ohhhh I did it the opposite way by subtracting 6x^2 from both sides

OpenStudy (anonymous):

Doesn't make any difference, you will still get the correct answer.

Vocaloid (vocaloid):

that should work too, you must have made an algebra mistake somewhere

OpenStudy (anonymous):

So x^2-2x-4 = 0?

OpenStudy (anonymous):

Should be +4

OpenStudy (anonymous):

Then just apply the equation?

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Oh right sorry +4

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

I got to here but I'm a little stuck.